Crude oil has sold at $21.75 a barrel. A barrel has a volume of 42 gallons.

What is the price of crude oil in ¢ /cup? (One gallon = 4 quarts; 4 cups = 1

quart)( ¢ =cents)

Group of answer choices





1
Expert's answer
2021-09-10T08:00:36-0400

We have to use the conversion factors to calculate the price in cents per cup:


"\text{Price: } P=\cfrac{\$21.75}{1\,barrel} \bigg( \dfrac{ 100\,¢}{\$\,1.00} \bigg) \bigg( \dfrac{1\,barrel}{42\,gallons} \bigg) \bigg( \dfrac{1\,gallon}{4\,quarts} \bigg) \bigg( \dfrac{1\,quart}{4\,cups} \bigg)
\\ \text{after substitution we find:}
\\ \implies \text{Price per cup: } P=\cfrac{3.24\,¢}{cup}"


In conclusion, the price of crude oil in ¢ /cup is 3.24 ¢ / cup.


Reference

  • Chang, R., & Goldsby, K. A. (2010). Chemistry. Chemistry, 10th ed.; McGraw-Hill Education: New York, NY, USA.